Determine the downward pull (due to surface tension) on a solid glass rod of diameter 3 cm when it stands vertically with its lower end submerged in water. Surface tension of water is 0.0735 N/m.
1
Expert's answer
2020-11-27T10:09:05-0500

Find the length of the surface that will be pulled into water:


"l=2\pi r=2\cdot3.14\cdot(0.03/2)="

The downward pull due to surface tension will be


"F=\sigma l=2\sigma\pi r=\\=2\cdot0.0735\cdot3.14\cdot0.015=6.93\cdot10^{-3}\text{ N}."